# BioCLIP 2: Emergent Properties from Scaling Hierarchical Contrastive Learning

### v1

@@ Abstract
Foundation models trained at scale exhibit remarkable emergent behaviors, learning new capabilities beyond their initial training objectives. We find such emergent behaviors in biological vision models via large-scale contrastive vision-language training. To achieve this, we first curate TreeOfLife-200M, comprising million images of living organisms, the largest and most diverse biological organism image dataset to date. We then train BioCLIP 2 on TreeOfLife-200M to distinguish different species. Despite the narrow training objective, BioCLIP 2 yields extraordinary accuracy when applied to various biological visual tasks such as habitat classification and trait prediction. We identify emergent properties in the learned embedding space of BioCLIP 2. At the inter-species level, the embedding distribution of different species aligns closely with functional and ecological meanings (e.g. beak sizes and habitats). At the intra-species level, instead of being diminished, the intra-species variations (e.g. life stages and sexes) are preserved and better separated in subspaces orthogonal to inter-species distinctions. We provide formal proof and analyses to explain why hierarchical supervision and contrastive objectives encourage these emergent properties. Crucially, our results reveal that these properties become increasingly significant with larger-scale training data, leading to a biologically meaningful embedding space.
@@ 1 Introduction
To probe these capabilities, we evaluate BioCLIP 2 on a variety of existing biological visual tasks, including habitat classification [16], trait prediction [17, 18], new-species identification [19], and agricultural disease detection [20]. These applications push beyond simple species recognition and apply directly to biodiversity conservation, trait organization, and agricultural health. Despite being trained primarily with species-level supervision, BioCLIP 2 outperforms both vision-language (e.g. SigLIP [21]) and vision-only baselines (e.g. DINOv2 [22]) by an average margin of on these tasks. We then look deeper into the embedding space of BioCLIP 2 and identify two emergent properties as the training scales up.
@@
At the intra-species level, contrary to the intuition that fine-grained differences collapse after extensive training [24, 25], BioCLIP 2 keeps the intra-species variations (e.g. life stages and sexes) distinct. On the right side of Figure 1, three species form tight clusters when projected onto the “species plane,” while their intra-species variations fan out along axes orthogonal to the plane. Such variation cues are not encoded in taxonomic labels. We theoretically prove that when species prototypes are nearly orthogonal (i.e., species are well separated), the contrastive objective prioritizes orthogonality between intra-species variations and inter-species differences over raw magnitude (Theorem 5.1). Furthermore, these variations are observed to be increasingly separable as training data scales up. This microstructure preserves the intra-species representational diversity without interfering with inter-species distinctions, enabling various attribute recognition applications (§5.1).
@@ 2 Related Work
Emergent properties in foundation models. Emergent properties refer to the capabilities implicitly acquired from the training process and generalized beyond the initial training objective. Large language models (LLMs) illustrate a variety of in-context learning skills after next-token-prediction pre-training [9, 26]. Such emergence also arises in computer vision. DINO learns semantic segmentation through purely visual self-supervision [27], whereas GroupViT learns semantic segmentation solely through text supervision [28]. The studies closest to this work are [29, 30]. Alper and Averbuch-Elor explore the visual-semantic hierarchies in CLIP models [29]. Although hierarchical semantic structures are never explicitly presented as the training supervision, CLIP models acquire the capability of matching images with varying levels of descriptions. Abbasi et al. discover that CLIP models possess disentangled representational sub-spaces for different factors of variations [30]. It allows CLIP models to generalize across compositional out-of-distribution concepts. This work leverages CLIP to distinguish different species with hierarchical labels, which is a different scenario from the above work. We investigate the emergent properties under this setting.
Computer vision for ecology & evolutionary biology. Ecology and evolutionary biology is a naturally challenging regime for computer vision systems due to long-tail distributions, extremely fine-grained classifications, and a wide variety of image distributions. Existing work formalizes these challenges into specific visual tasks such as attribute prediction [NeWT, 18], trait prediction [FishNet, 16], and plant disease detection [PlantDoc, 20].
Recent advancements in computer vision have led to the development of foundation models for biological applications. BioCLIP incorporates taxonomic labels in the vision-language contrastive training, yielding promising species classification accuracy [13]. Follow-up work scaled data to M images [BioTrove, 14], specialized the data to camera traps [CATALOG and WildCLIP, 31, 32], and added additional model modalities [TaxaBind, 33]. We investigate both data and model scaling, with a focus on both broad biological applications and any emergent properties after extensive training.
@@ 3.2 Data Curation & Filtering
We retrieve data from data providers using distributed-downloader. The initial retrieval gives us images and unique taxonomic hierarchies. We then cleaned the data, focusing on (1) aligning taxonomic labels, (2) image quality, and (3) eliminating duplication and data leakage. We summarize these efforts here, with more details provided in §I.
@@ 4.1 Species Classification Performance
We evaluate BioCLIP 2 on species classification tasks in Table 1. We use the same benchmarks as BioCLIP [13], including seven tasks from Meta-Album [56] and Rare Species [57]. We substitute Birds-525 [58] with NABirds [59] due to data inaccessibility. Additionally, we collect a test set of camera trap images from the Labeled Information Library of Alexandria: Biology and Conservation (LILA-BC) [60, 61, 62, 63, 64, 65] to illustrate more realistic species classification applications in the wild. The results suggest substantial improvements of BioCLIP 2 over BioCLIP. Particularly, attributed to more comprehensive species and image type coverage of TreeOfLife-200M, we observe over zero-shot improvement on CameraTrap, Fungi, and Rare Species. On average, BioCLIP 2 surpasses the second-best model by and provides a improvement over the original CLIP model that serves as weight initialization. Information of baselines is in §E.
@@ 5.1 Beyond Species Classification
Biology’s organization extends beyond species taxonomies; if scaling truly induces emergent behavior, model representations learned through species-level supervision should transfer to problems far removed from species classification. We collect and benchmark models on five visual benchmarks that push past species ID: habitat classification (ecological context) [16], trait prediction (evolutionary studies) [18, 17], new-species identification (biodiversity monitoring) [19], and agricultural disease detection [20]. For each task, we keep the evaluated models frozen and extract the corresponding sample embeddings. The embeddings are subsequently processed using machine-learning techniques (e.g. support vector classifiers). Detailed evaluation procedures are listed in §H.
Table 2 presents the performance comparison among BioCLIP 2, vision-language baselines, and vision-only models. Although no information on these tasks is explicitly described during training, BioCLIP 2 yields an average performance improvement of over the original CLIP baseline. DINOv2 is commonly believed to capture fine-grained visual features and is adopted for diverse visual tasks [22, 66, 67]. Nevertheless, BioCLIP 2 yields a performance gap over DINOv2.
@@ 5.2 Scaling Trends
Next, we investigate how scale affects representations within species. We collect two groups of images with intra-species appearance variations: life stage variations from NeWT [18] and sex variations from NABirds [59]. We ask whether scaling hierarchical contrastive training collapses all images of one species onto a single prototype or still distinguishes juveniles from adults and males from females. We accordingly design two complementary tasks for each type of variation: (i) alignment, where a species classifier trained on one variant (e.g. juvenile images) is expected to recognize the species on the other variant (e.g. adult images), and (ii) differentiation, where the task is to tell the variants apart (e.g. juvenile vs. adult). As illustrated in 3(b), data scale steadily improves cross-variant species recognition. But at the same time, the model also becomes better at distinguishing the variants themselves.
@@ 5.4 Formal Analysis
Furthermore, we look at the separation between these variations using the Fisher Discriminant Ratio (FDR) metric [71]. Given two variant classes and , FDR is defined by their embedding mean and standard deviation : . The increasing trends in 3(c) support that in addition to orthogonality, the intra-species variations are more separable as the training scales up. This trend is also qualitatively evidenced by the top row of Figure 5. These theoretical and empirical insights validate that BioCLIP 2 learns to preserve and separate intra-species variations without explicit training constraints. We provide more qualitative analyses in §F.2.
@@ 7 Conclusion
We are grateful to Kakani Katija and Dirk Steinke for helpful conversations regarding use and integration of FathomNet and BIOSCAN-5M, respectively, as well as Steven Formel and Markus Döring for GBIF. We thank Marie Grosjean for comparative methods for filtering citizen science images and Dylan Verheul for assistance with acquiring images from observation.org from GBIF. We thank Suren Byna for a helpful conversation on early dataset design decisions. We thank Doug Johnson for his collaboration in hosting this large dataset on the Ohio Supercomputer Center research storage file system.
@@
In §J we list the contribution of each author.
@@ Appendix A Limitations
Further emphasis on the impact of citizen science images in amassing larger representations of species: the most prevalent taxonomic classes, flowering plants, insects, birds, mushrooms, and mammals, have millions of representative images, while the least-represented, microscopic organisms (e.g., bacteria, viruses), have a dozen or less.
@@ Appendix B Broader Impacts
BioCLIP 2 and TreeOfLife-200M provide great potential to improve and enhance existing conservation efforts, in particular by facilitating recognition of threatened species. As noted in §3.3, TreeOfLife-200M has expansive coverage of threatened species, as classified by IUCN. It additionally builds on the coverage of species considered to be Data Deficient. Based on the emergent properties displayed by BioCLIP 2, there is potential to add in the effort to understand the risks facing these species that cannot currently be classified by IUCN due to lack of available information, as suggested in [74]. These designations are crucial to the international effort to protect biodiversity across the planet.
Unfortunately, as with many open-source efforts to further conservation goals, there is potential for bad actors to make use of these tools for malicious purposes. Though the improvement on threatened species could make it easier for poachers to identify protected species, these types of tools are a force-multiplier to monitor illicit trade and sales of these same species. The primary risk to endangered species comes from disclosure of precise location information rather than improved classification capability [75]. Our data does not provide geo-tagged information of the organisms included, minimizing the vulnerabilities that could be used in poaching.
@@
Define . Then collecting first order terms around results in the following:
@@
The Hessian matrix lies in the span of species as each term is an outer product of a prototype. Therefore, as long as the residual is distributed in a subspace orthogonal to the species differences, the scale of will not interfere with the overall contrastive optimization. ∎
@@ Appendix D Training Implementation Details
We list the adopted hyper-parameters in Table 4 and Table 4 for the BioCLIP 2 training and ablation study, respectively. TreeOfLife-10M is used for the ablation study. The batch size presented in both tables is the size per GPU. In addition to the larger GPU number and smaller batch size compared with the training in BioCLIP due to the larger model size, another important modification is the introduction of experience replay. An additional visual projector is introduced as described in §F.1. Beyond the visual projector, the model architecture is kept the same as that of CLIP [12]. During the evaluation, all the embeddings are extracted with the visual projector for hierarchical label matching to avoid extra influence.
@@ Appendix E Baseline Details
DINOv2. Besides supervised pre-training, we also evaluate the performance of DINOv2, which is pre-trained in an unsupervised way [22]. The backbone architecture is ViT-large and the patch size is . The model can be downloaded from Hugging Face. Similarly, we only run few-shot and non-species classification evaluations with DINOv2.
@@ F.1 CLIP Training Data Replay
We evaluate different replay settings quantitatively in Table 5. The “No Replay” row shows the baseline performance applying biological contrastive training on top of the pre-trained CLIP model, where a performance improvement is achieved. When the replay data is added and share the visual projector with biological images, we observe conflicted performance change. On Herb. 19 [19] and PlantDoc [20], more than improvement is acquired. However, the species classification accuracy on Rare Species drops by . We attribute the inconsistent performance change to the distribution gap between biological and replay text labels. After applying a separate visual projector for the replay data, we observe overall improved performance across multiple benchmarks. At the same time, we also notice that replay has limited influence on benchmarks like AwA2 [17]. Additionally, we evaluate the preservation of general knowledge understanding on INQUIRE-Rerank [55]. Applying contrastive training with taxonomic labels slightly hurts the performance, while the single-projector replay fails to retain the understanding. Comparatively, the adopted separate-projector replay improves the performance by . Given that there is still a distribution gap between the taxonomic labels and natural language, we do not treat INQUIRE as one of our main focuses in this work.
@@ F.2 Qualitative Analyses
Embedding distribution of Darwin’s finches. In Figure 1 we show that the embedding distribution of Darwin’s finches aligns with the beak size. Here we further visualize the distribution under different training scales in Figure 8. Based on the genome-based phylogeny, warbler finches were the most ancient branches, while tree finches and ground finches form the recent branches [23]. Among these species, warbler finches have the smallest beak, convenient for extracting tiny arthropods from leaves. Comparatively, ground finches have larger beaks, which are more suitable for cracking seeds and nuts. In the original CLIP embedding space, the warbler finches and tree finches are mixed. As the training data scales up, these two groups are separated, and the relative geometric relationship of all the finches aligns with their phylogeny tree. While the species separation is induced by taxonomic supervision, BioCLIP 2’s embedding space again illustrates emergent higher-level biological meaningfulness after extensive training.
Embedding distribution of Sex data. Similar to Figure 5, we visualize the 2D and 3D distributions of embeddings from 3 species of the Sex data in Figure 9. We can draw conclusions similar to that obtained from Figure 5. When no training data is incorporated, the embeddings extracted by the original CLIP visual encoder (leftmost sub-figure) demonstrate large portions of overlap between male and female images. After extensive vision-language contrastive training, the embeddings present clear decision boundaries between the two variants. Furthermore, as evidenced in the 3D distribution, the embeddings within each species form more compact clusters when projected onto the species span (the gray plane). Comparatively, instead of being eliminated after contrastive training, the intra-species variations are embedded in the subspace orthogonal to the inter-species differences. The extensive training facilitates BioCLIP 2 to acquire a biologically meaningful embedding space, highlighting its value in serving as a biology foundation model.
@@
We observe that FDR between two intra-species variation classes is increasing as the training data scales up. More specifically, we look into the numerator term (difference between two class centers) and the denominator term (the variation of features). We visualize the curve of the numerator and denominator terms in 11(a) and 11(b) for life stage data and sex data, respectively. The y-axes are scaled to the same maximum ratio to the minimum value. As the number of training samples scales up from M to M, the denominator term goes through a larger decay than the numerator term. The numerical results further support the increasing FDR trend.

### v2

@@ Abstract
Foundation models trained at scale exhibit remarkable emergent behaviors, learning new capabilities beyond their initial training objectives. We find such emergent behaviors in biological vision models via large-scale contrastive vision-language training. To achieve this, we first curate TreeOfLife-200M, comprising million images of living organisms, the largest and most diverse biological organism image dataset to date. We then train BioCLIP 2 on TreeOfLife-200M to distinguish different species. Despite the narrow training objective, BioCLIP 2 yields extraordinary accuracy when applied to various biological visual tasks such as habitat classification and trait prediction. We identify emergent properties in the learned embedding space of BioCLIP 2. At the inter-species level, the embedding distribution of different species aligns closely with functional and ecological meanings (e.g., beak sizes and habitats). At the intra-species level, instead of being diminished, the intra-species variations (e.g., life stages and sexes) are preserved and better separated in subspaces orthogonal to inter-species distinctions. We provide formal proof and analyses to explain why hierarchical supervision and contrastive objectives encourage these emergent properties. Crucially, our results reveal that these properties become increasingly significant with larger-scale training data, leading to a biologically meaningful embedding space.
@@ 1 Introduction
To probe these capabilities, we evaluate BioCLIP 2 on a variety of existing biological visual tasks, including habitat classification [16], trait prediction [17, 18], new-species identification [19], and agricultural disease detection [20]. These applications push beyond simple species recognition and apply directly to biodiversity conservation, trait organization, and agricultural health. Despite being trained primarily with species-level supervision, BioCLIP 2 outperforms both vision-language (e.g., SigLIP [21]) and vision-only baselines (e.g., DINOv3 [22]) by an average margin of on these tasks. We then look deeper into the embedding space of BioCLIP 2 and identify two emergent properties as the training scales up.
@@
At the intra-species level, contrary to the intuition that fine-grained differences collapse after extensive training [24, 25], BioCLIP 2 keeps the intra-species variations (e.g., life stages and sexes) distinct. On the right side of Figure 1, three species form tight clusters when projected onto the “species plane,” while their intra-species variations fan out along axes orthogonal to the plane. Such variation cues are not encoded in taxonomic labels. We theoretically prove that when species prototypes are nearly orthogonal (i.e., species are well separated), the contrastive objective prioritizes orthogonality between intra-species variations and inter-species differences over raw magnitude (Theorem 5.1). Furthermore, these variations are observed to be increasingly separable as training data scales up. This microstructure preserves the intra-species representational diversity without interfering with inter-species distinctions, enabling various attribute recognition applications (§5.1).
@@ 2 Related Work
Emergent properties in foundation models. Emergent properties refer to the capabilities implicitly acquired from the training process and generalized beyond the initial training objective. Large language models (LLMs) illustrate a variety of in-context learning skills after next-token-prediction pre-training [9, 26]. Such emergence also arises in computer vision [27]. DINO learns semantic segmentation through purely visual self-supervision [28], whereas GroupViT learns semantic segmentation solely through text supervision [29]. The studies closest to this work are [30, 31]. Alper and Averbuch-Elor explore the visual-semantic hierarchies in CLIP models [30]. Although hierarchical semantic structures are never explicitly presented as the training supervision, CLIP models acquire the capability of matching images with varying levels of descriptions. Abbasi et al. discover that CLIP models possess disentangled representational sub-spaces for different factors of variations [31]. It allows CLIP models to generalize across compositional out-of-distribution concepts. This work leverages CLIP to distinguish different species with hierarchical labels, which is a different scenario from the above work. We investigate the emergent properties under this setting.
Computer vision for ecology & evolutionary biology. Ecology and evolutionary biology are naturally challenging for computer vision systems due to long-tail distributions, extremely fine-grained classifications, and a wide variety of image distributions. Existing work formalizes these challenges into specific visual tasks such as attribute prediction [NeWT, 18], trait prediction [FishNet, 16], and plant disease detection [PlantDoc, 20].
Recent advancements in computer vision have led to the development of foundation models for biological applications. BioCLIP incorporates taxonomic labels in the vision-language contrastive training, yielding promising species classification accuracy [13]. Follow-up work scaled data to M images [BioTrove, 14], specialized the data to camera traps [CATALOG and WildCLIP, 32, 33], and added additional model modalities [TaxaBind, 34]. We investigate both data and model scaling, with a focus on both broad biological applications and any emergent properties after extensive training.
@@ 3.2 Data Curation & Filtering
We retrieve data from data providers using distributed-downloader [45]. The initial retrieval gives us images and unique taxonomic hierarchies. We then cleaned the data, focusing on (1) aligning taxonomic labels, (2) image quality, and (3) eliminating duplication and data leakage. We summarize these efforts here, with more details provided in §I.
@@ 4.1 Species Classification Performance
We evaluate BioCLIP 2 on species classification tasks in Table 1. We use the same benchmarks as BioCLIP [13], including seven tasks from Meta-Album [58] and Rare Species [59]. We substitute Birds-525 [60] with NABirds [61] due to data inaccessibility. Additionally, we collect a test set of IDLE-OO Camera Traps from the Labeled Information Library of Alexandria: Biology and Conservation (LILA-BC) [62, 63, 64, 65, 66, 67] to illustrate more realistic species classification applications in the wild. The results suggest substantial improvements of BioCLIP 2 over BioCLIP. Particularly, attributed to more comprehensive species and image type coverage of TreeOfLife-200M, we observe over zero-shot improvement on Camera Trap, Fungi, and Rare Species. On average, BioCLIP 2 surpasses the second-best model by and provides a improvement over the original CLIP model that serves as weight initialization. Information on baselines is in §E.
@@ 5.1 Beyond Species Classification
Biology’s organization extends beyond species taxonomies; if scaling truly induces emergent behavior, model representations learned through species-level supervision should transfer to problems far removed from species classification. We collect and benchmark models on five visual benchmarks that push past species ID: habitat classification (ecological context) [16], trait prediction (evolutionary studies) [18, 17], new-species identification (biodiversity monitoring) [19], and agricultural disease detection [20]. For each task, we keep the evaluated models frozen and extract the corresponding sample embeddings. The embeddings are subsequently processed using machine-learning techniques (e.g., support vector classifiers). Detailed evaluation procedures are listed in §H.
Table 2 presents the performance comparison among BioCLIP 2, vision-language baselines, and vision-only models. Although no information on these tasks is explicitly described during training, BioCLIP 2 yields an average performance improvement of over the original CLIP baseline. DINOv3 is commonly believed to capture fine-grained visual features and is adopted for diverse visual tasks [22, 68, 69]. Nevertheless, BioCLIP 2 yields an performance gap over DINOv3.
@@ 5.2 Scaling Trends
Next, we investigate how scale affects representations within species. We collect two groups of images with intra-species appearance variations: life stage variations from NeWT [18] and sex variations from NABirds [61]. We ask whether scaling hierarchical contrastive training collapses all images of one species onto a single prototype or still distinguishes juveniles from adults and males from females. We accordingly design two complementary tasks for each type of variation: (i) alignment, where a species classifier trained on one variant (e.g., juvenile images) is expected to recognize the species on the other variant (e.g., adult images), and (ii) differentiation, where the task is to tell the variants apart (e.g., juvenile vs. adult). As illustrated in 3(b), data scale steadily improves cross-variant species recognition. But at the same time, the model also becomes better at distinguishing the variants themselves.
@@ 5.4 Formal Analysis
Furthermore, we look at the separation between these variations using the Fisher Discriminant Ratio (FDR) metric [73]. Given two variant classes and , FDR is defined by their embedding mean and standard deviation : . The increasing trends in 3(c) support that, in addition to orthogonality, the intra-species variations are more separable as the training scales up. This trend is also qualitatively evidenced by the top row of Figure 5. These theoretical and empirical insights validate that BioCLIP 2 learns to preserve and separate intra-species variations without explicit training constraints. We provide more qualitative analyses in §F.2.
@@ 7 Conclusion
We are grateful to Kakani Katija and Dirk Steinke for helpful conversations regarding use and integration of FathomNet and BIOSCAN-5M, respectively, as well as Stephen Formel and Markus Döring for GBIF. We thank Marie Grosjean for comparative methods for filtering citizen science images and Dylan Verheul for assistance with acquiring images from observation.org from GBIF. We thank Suren Byna for a helpful conversation on early dataset design decisions. We thank Doug Johnson for his collaboration in hosting this large dataset on the Ohio Supercomputer Center research storage file system.
@@
In §LABEL:sec:appendix-contribution we list the contribution of each author.
@@ Appendix A Limitations
Further emphasis on the impact of citizen science images in amassing larger representations of species: the most prevalent taxonomic classes, flowering plants, insects, birds, mushrooms, and mammals, have millions of representative images, while the least-represented, microscopic organisms (e.g., bacteria, viruses), have a dozen or fewer.
@@ Appendix B Broader Impacts
BioCLIP 2 and TreeOfLife-200M provide great potential to improve and enhance existing conservation efforts, in particular by facilitating recognition of threatened species. As noted in §3.3, TreeOfLife-200M has expansive coverage of threatened species, as classified by IUCN. It additionally builds on the coverage of species considered to be Data Deficient. Based on the emergent properties displayed by BioCLIP 2, there is potential to add to the effort to understand the risks facing these species that cannot currently be classified by IUCN due to lack of available information, as suggested in [76]. These designations are crucial to the international effort to protect biodiversity across the planet.
Unfortunately, as with many open-source efforts to further conservation goals, there is potential for bad actors to make use of these tools for malicious purposes. Though the improvement on threatened species could make it easier for poachers to identify protected species, these types of tools are a force-multiplier to monitor illicit trade and sales of these same species. The primary risk to endangered species comes from disclosure of precise location information rather than improved classification capability [77]. Our data does not provide geo-tagged information on the organisms included, minimizing the vulnerabilities that could be used in poaching.
@@
Define . Then collecting first-order terms around results in the following:
@@
The Hessian matrix lies in the span of species, as each term is an outer product of a prototype. Therefore, as long as the residual is distributed in a subspace orthogonal to the species differences, the scale of will not interfere with the overall contrastive optimization. ∎
@@ Appendix D Training Implementation Details
We list the adopted hyperparameters in Table 4 and Table 4 for the BioCLIP 2 training and ablation study, respectively. TreeOfLife-10M is used for the ablation study. The batch size presented in both tables is the size per GPU. In addition to the larger GPU number and smaller batch size compared with the training in BioCLIP due to the larger model size, another important modification is the introduction of experience replay. An additional visual projector is introduced as described in §F.1. Beyond the visual projector, the model architecture is kept the same as that of CLIP [12]. During the evaluation, all the embeddings are extracted with the visual projector for hierarchical label matching to avoid extra influence.
@@ Appendix E Baseline Details
DINOv3. Besides supervised pre-training, we also evaluate the performance of DINOv3, which is pre-trained in an unsupervised way [22]. The backbone architecture is ViT-large, and the patch size is . The model can be downloaded from Hugging Face. Similarly, we only run few-shot and non-species classification evaluations with DINOv3.
@@ F.1 CLIP Training Data Replay
We evaluate different replay settings quantitatively in Table 5. The “No Replay” row shows the baseline performance applying biological contrastive training on top of the pre-trained CLIP model, where a performance improvement is achieved. When the replay data is added, which shares the visual projector with biological images, we observe a conflicting performance change. On Herb. 19 [19] and PlantDoc [20], more than improvement is acquired. However, the species classification accuracy on Rare Species drops by . We attribute the inconsistent performance change to the distribution gap between biological and replay text labels. After applying a separate visual projector for the replay data, we observe overall improved performance across multiple benchmarks. At the same time, we also notice that replay has limited influence on benchmarks like AwA2 [17]. Additionally, we evaluate the preservation of general knowledge understanding on INQUIRE-Rerank [57]. Applying contrastive training with taxonomic labels slightly hurts the performance, while the single-projector replay fails to retain the understanding. Comparatively, the adopted separate-projector replay improves the performance by . Given that there is still a distribution gap between the taxonomic labels and natural language, we do not treat INQUIRE as one of our main focuses in this work.
@@ F.2 Qualitative Analyses
Embedding distribution of Darwin’s finches. In Figure 1, we show that the embedding distribution of Darwin’s finches aligns with the beak size. Here we further visualize the distribution under different training scales in Figure 8. Based on the genome-based phylogeny, warbler finches were the most ancient branches, while tree finches and ground finches form the recent branches [23]. Among these species, warbler finches have the smallest beak, convenient for extracting tiny arthropods from leaves. Comparatively, ground finches have larger beaks, which are more suitable for cracking seeds and nuts. In the original CLIP embedding space, the warbler finches and tree finches are mixed. As the training data scales up, these two groups are separated, and the relative geometric relationship of all the finches aligns with their phylogeny tree. While the species separation is induced by taxonomic supervision, BioCLIP 2’s embedding space again illustrates emergent higher-level biological meaningfulness after extensive training.
Embedding distribution of Sex data. Similar to Figure 5, we visualize the 2D and 3D distributions of embeddings from 3 species of the Sex data in Figure 9. We can draw conclusions similar to those obtained from Figure 5. When no training data is incorporated, the embeddings extracted by the original CLIP visual encoder (leftmost sub-figure) demonstrate large portions of overlap between male and female images. After extensive vision-language contrastive training, the embeddings present clear decision boundaries between the two variants. Furthermore, as evidenced in the 3D distribution, the embeddings within each species form more compact clusters when projected onto the species span (the gray plane). Comparatively, instead of being eliminated after contrastive training, the intra-species variations are embedded in the subspace orthogonal to the inter-species differences. The extensive training facilitates BioCLIP 2 to acquire a biologically meaningful embedding space, highlighting its value in serving as a biology foundation model.
@@
We observe that FDR between two intra-species variation classes is increasing as the training data scales up. More specifically, we look into the numerator term (the difference between two class centers) and the denominator term (the variation of features). We visualize the curve of the numerator and denominator terms in 11(a) and 11(b) for life stage data and sex data, respectively. The y-axes are scaled to the same maximum ratio to the minimum value. As the number of training samples scales up from M to M, the denominator term goes through a larger decay than the numerator term. The numerical results further support the increasing FDR trend.